# How much do contextualized representations encode long-range context?

### v1

@@ 6 Related Work
Contextualized representations have been shown to encode useful linguistic features (Liu et al., 2019; Hewitt and Manning, 2019) examined with linear probing. In contrast, this work analyzes the geometry of the representations built on previous analyses (Ethayarajh, 2019; Cai et al., 2021). Prior works often analyze sentence-level data, while our focus is the representations given long-range context. Perturbations are often employed to analyze the effect of context (Khandelwal et al., 2018; Sun et al., 2021) and the robustness in adversarial scenarios (Li et al., 2021); more controlled causal interventions are used to study other phenomena such as verbatim memorization in language models (Huang et al., 2024a).
Long-range context often touches the length extrapolation regime. Better length generalization (Lake and Baroni, 2017; Anil et al., 2022; Deletang et al., 2023; Zhou et al., 2023) performance is often driven by novel designs, such as modifications to position encoding (Chen et al., 2023; Sun et al., 2023b; Xiong et al., 2024; Jin et al., 2024), attention (Wu et al., 2022; Wang et al., 2023), adaptive layer depth (Fan et al., 2024) in Transformer and completely new architectures (Gu and Dao, 2024; Poli et al., 2023; Sun et al., 2023a; Peng et al., 2023; Ma et al., 2024; Yang et al., 2024), which deserve additional investigation. Many of the designs can be considered as applying “regularization” to the global long-range context, and are relevant to Rosenfeld (1996)’s max entropy approach mixing local and global context (Bau and Andreas, 2021). Recent novel designs are often evaluated using perplexity and downstream tasks. Recent work (Lu et al., 2024) demonstrates a linear relationship between perplexity and downstream tasks. Our results suggest a more intricate relationship between these two.
Anisotropic representations have been observed by prior works (Gao et al., 2019) and were shown to be alleviated by contrastive method (Su et al., 2022) and spectrum control (Wang et al., 2020). Recent work suggests anisotropy is inherent to self-attention (Godey et al., 2024) and whether is it harmful or not actually depends on the downstream task (Ait-Saada and Nadif, 2023). We study anisotropy as an intermediate step towards understanding contextualized representations, by relating anisotropy to an intrinsic property of the input string, inspired by the view of language modeling as compression (Deletang et al., 2024; Gu et al., 2024). Finally, our analysis suggests better contextualization of the entire long-range context in hybrid models, echoing recent positive results from hybridization of attention and other layer modules (Team, 2024; Ren et al., 2024; Waleffe et al., 2024).
@@ Appendix B Model hyperparameters & other details
The hyperparameters used by GPT+RoPE and GPT+ALiBi are shown in Table 2. For ALiBi, we use the default slopes specified in the NeMo (Kuchaiev et al., 2019) framework. [^6] https://github.com/NVIDIA/NeMo/blob/main/nemo/collections/nlp/modules/common/megatron/position_embedding/alibi_relative_position_embedding.py All architectures share the same vocab size and training data. For other architectures, we modify the model depth to achieve roughly same number of parameters ($\sim$0.47B) across all models. When re-implementing alternative architectures, we find it crucial to cast data into torch.float32 during context mixing (e.g. cumsum of forget gates in mLSTM) for stabilized training. For mLSTM we use conv1d kernel size 4, qkv number of heads 4, projection factor 2. For HybridMamba, we mix 8% attention layers with 62% of Mamba-2 layers and 30% MLP layers. For Griffin, we insert local attention layers (window size = 1024) every two layers of RG-LRU layers. For both hybrid models, the first layer is consistently a non-attention layer. We anneal to final learning rate of 0. The peak learning rate for Griffin and mLSTM is 0.001, with the rest of parameters the same as specified in Table 2.
@@ Appendix D Additional discussion on Anisotropy and sequence complexity
In Figure 8, we plot the retrieval accuracy of llama3.1-8b-base with increasing number of distractor needles in the “haystack” of highly compressible repetitive sentences. While it remains unclear in prior works whether increased anisotropy hurts downstream task performance or not, we show here that the increased anisotropy, which happens alongside the increase in prefix compression rate, co-occurs with decreased in-context retrieval performance.
@@
We have shown that as the prefix becomes less compressible, the contextualized representations become increasingly aligned in certain directions. But what do these representations collapse into? Our preliminary analysis identifies two distinct modes: (1) the representation corresponding to a uniform distribution over the vocabulary, and (2) the representation corresponding to the unigram prior of the pre-training corpus. Figure 9 display two cases where representations are anisotropic: (left) sequences that exceed the maximum training length, especially for GPT+RoPE without training-free extension, and (right) sequences that are less compressible due to the absence of patterns. Consider a sequence of length $L$ consisting of tokens from a finite vocabulary of size $|V|$. Increasing $L$ leads to exponential increase of high complexity strings for which models demonstrate decreased representational capacity, as reflected by the increase in anisotropy, or deterioration into predicting unigram prior. The reduced isotropy can potentially explain common observations of repetitive output of frequent words (e.g. “aaaaa”, “the the”) given long prefixes, while more evidence is needed. Precisely quantifying the “regularization” required for the expanding prefix string can potentially help guide better design of new language models.

### v2

@@ 6 Related Work
Contextualized representations have been shown to encode useful linguistic features (Liu et al., 2019; Hewitt and Manning, 2019) examined with linear probing. Recent work (Morris et al., 2023) studies these representations by inverting them back into short sequences. In contrast, this work focuses on long inputs and employs a different methodology by analyzing the representation geometry, building on prior analyses (Ethayarajh, 2019; Cai et al., 2021). Perturbations are often employed to analyze the effect of context (Khandelwal et al., 2018; Sun et al., 2021) and the robustness in adversarial scenarios (Li et al., 2021); more controlled causal interventions are used to study other phenomena such as verbatim memorization in language models (Huang et al., 2024a).
Long-range context often touches the length extrapolation regime. Better length generalization (Lake and Baroni, 2017; Anil et al., 2022; Deletang et al., 2023; Zhou et al., 2023) performance is often driven by novel designs, such as modifications to position encoding (Chen et al., 2023; Sun et al., 2023b; Xiong et al., 2024; Jin et al., 2024), attention (Wu et al., 2022; Wang et al., 2023), adaptive layer depth (Fan et al., 2024) in Transformer and completely new architectures (Gu and Dao, 2024; Poli et al., 2023; Sun et al., 2023a; Peng et al., 2023; Ma et al., 2024; Yang et al., 2024), which deserve additional investigation. Many of the designs can be considered as applying “regularization” to the global long-range context, and are relevant to Rosenfeld (1996)’s max entropy approach mixing local and global context (Bau and Andreas, 2021). Recent novel designs are often evaluated using perplexity and downstream tasks. Recent work (Lu et al., 2024) demonstrates a linear relationship between perplexity and long-context downstream tasks. Our results suggest a more intricate relationship between these two.
Anisotropic representations have been observed by prior works (Gao et al., 2019) and were shown to be alleviated by contrastive method (Su et al., 2022; Gao et al., 2021), spectrum control (Wang et al., 2020), and proper regularization (Zhang et al., 2020). Recent work suggests that anisotropy is inherent to self-attention (Godey et al., 2024), and it is greatly affected by token frequency (Zhou et al., 2021; Puccetti et al., 2022). While anisotropic representations are often considered harmful, recent work shows it depends on the downstream task (Ait-Saada and Nadif, 2023). We study anisotropy as an intermediate step towards understanding contextualized representations, by relating anisotropy to an intrinsic property of the input string, inspired by the view of language modeling as compression (Deletang et al., 2024; Gu et al., 2024). Finally, our analysis suggests better contextualization of the entire long-range context in hybrid models, echoing recent positive results from hybridization of attention and other layer modules (Team, 2024; Ren et al., 2024; Waleffe et al., 2024).
@@ Appendix B Model hyperparameters & other details
The hyperparameters used by GPT+RoPE and GPT+ALiBi are shown in Table 2. For ALiBi, we use the default slopes specified in the NeMo (Kuchaiev et al., 2019) framework. [^6] https://github.com/NVIDIA/NeMo/blob/main/nemo/collections/nlp/modules/common/megatron/position_embedding/alibi_relative_position_embedding.py All architectures share the same vocab size and training data. The models we pre-trained do not tie the input and output embeddings. For other architectures, we modify the model depth to achieve roughly same number of parameters ($\sim$0.47B) across all models. When re-implementing alternative architectures, we find it crucial to cast data into torch.float32 during context mixing (e.g. cumsum of forget gates in mLSTM) for stabilized training. For mLSTM we use conv1d kernel size 4, qkv number of heads 4, projection factor 2. For HybridMamba, we mix 8% attention layers with 62% of Mamba-2 layers and 30% MLP layers. For Griffin, we insert local attention layers (window size = 1024) every two layers of RG-LRU layers. For both hybrid models, the first layer is consistently a non-attention layer. We anneal to final learning rate of 0. The peak learning rate for Griffin and mLSTM is 0.001, with the rest of parameters the same as specified in Table 2.
@@ Appendix D Additional discussion on Anisotropy and sequence complexity
In Figure 8, we plot the retrieval accuracy of llama3.1-8b-base with increasing number of distractor needles in the “haystack” of highly compressible repetitive sentences. While it remains unclear in prior works whether increased anisotropy hurts downstream task performance or not, we show here that the increased anisotropy, which happens alongside the increase in prefix compression rate, co-occurs with decreased in-context retrieval performance. The decreased downstream task performance is also consistent with prior work (Machina and Mercer, 2024).
@@
We have shown that as the prefix becomes less compressible, the contextualized representations become increasingly aligned in certain directions. But what do these representations collapse into? Our preliminary analysis identifies two distinct modes: (1) the representation corresponding to a uniform distribution over the vocabulary, and (2) the representation corresponding to the unigram prior of the pre-training corpus. Figure 9 displays two cases where representations are anisotropic: (left) sequences that exceed the maximum training length, especially for GPT+RoPE without training-free extension, and (right) sequences that are less compressible due to the absence of patterns. Consider a sequence of length $L$ consisting of tokens from a finite vocabulary of size $|V|$. Increasing $L$ leads to exponential increase of high complexity strings for which models demonstrate decreased representational capacity, as reflected by the increase in anisotropy, or deterioration into predicting unigram prior. The reduced isotropy can potentially explain common observations of repetitive output of frequent words (e.g. “aaaaa”, “the the”) given long prefixes, while more evidence is needed. Precisely quantifying the “regularization” required for the expanding prefix string can potentially help guide better design of new language models.